OpenStudy (anonymous):

evaluate this integral ∫x^2 +4 /x^2 -4 dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?} x ^{2}dx+\int\limits_{?}^{?} 4dx/x ^{2}-\int\limits_{?}^{?} 4dx\]